OpenStudy (anonymous):

7. The hypotenuse of a right triangle is 37. The longer leg is one less than twice the shorter leg. Set up a solve a Pythagorean Theorem solution to find the lengths of the legs of the triangle

OpenStudy (anonymous):

I have let a=x and b=2x-1

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and let c =37

OpenStudy (anonymous):

(x)^2-(2x-1)^2=37

OpenStudy (anonymous):

dont we sqaure 37 too?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

which is friggin' huge= 1369

OpenStudy (anonymous):

so I have 5x^2+4x-1=1369 ?

OpenStudy (anonymous):

but I can'tget past that. I know I need -b +- square root of b^2-4ac all over 2a but it's just such a hunge number o.o

OpenStudy (anonymous):

\[x^2+(2x-1)^2=37^2\]

OpenStudy (anonymous):

it should be -4x in that equation you set up

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

but I do \[4+-\sqrt{16-4(5)(-1370)}/10\] ?

OpenStudy (anonymous):

it's just that number is mega huge, and it doesn't seem right?

OpenStudy (anonymous):

I know the answers are 17 & 33, but I can't figure out how to get there

OpenStudy (anonymous):

i got it, give me a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you sure that the problem is written correctlly

OpenStudy (anonymous):

can you double check that the problem is written correctly?

OpenStudy (anonymous):

im sorry yes

OpenStudy (anonymous):

7. The hypotenuse of a right triangle is 37. The longer leg is one less than twice the shorter leg. Set up a solve a Pythagorean Theorem solution to find the lengths of the legs of the triangle. ans: 17 and 33

OpenStudy (anonymous):

that is how it is written. there was one question before that no matter what I just couldn't get the same answer. I wonder if the teacher had some wrong answers on there?

OpenStudy (anonymous):

what if it was 36 instead of 37?

OpenStudy (anonymous):

or even just 6?

OpenStudy (anonymous):

or 7?

OpenStudy (anonymous):

i think something may be off

OpenStudy (anonymous):

its just weird cause the it gave such a weird quadratic to solve. Usually these problems are a piece of cake, very routine. I assume your in highschool or something?

OpenStudy (anonymous):

no college

OpenStudy (anonymous):

right, either getting ready to leave highschool or in college i meant

OpenStudy (anonymous):

yeah, but still a problem like this should be routine

OpenStudy (anonymous):

but 37 squared is just so huge, it really doesn't make sense. I'm just going to ask him on tuesday. since all the examples use 9 or 6 or some single didgt...

OpenStudy (anonymous):

I did give you a good answer through for digging through it with me. It just doesn't make sense at all.

OpenStudy (anonymous):

yeah, thanks, its just 37 is a weird number for hypothense, and the quadratic is also not making sense, it should factor easily

OpenStudy (anonymous):

well we're suing the quadratic forumla, but still 37^2 is SOO huge! and then to make it a negitice just doesn't work.

OpenStudy (anonymous):

using* negitive*

OpenStudy (anonymous):

anyways, thank you for trying to help. <3